NEW ANALYSIS OF THE ν_{3} BAND OF HDCO (MONODEUTERATED FORMALDEHYDE) IN THE 5.8 μm REGION

L. GOMEZ, LAboratoire de Dynamique, Interactions et Réactivité (LADIR, CNRS UMR 7075) Université Pierre et Marie Curie - Paris VI ; Case courrier 49, Bât F 74, 4, place Jussieu, 75252 Paris Cedex, France.; \underline{A.PERRIN}, Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS, Université Paris 12, 61 Av du General de Gaulle, 94010 Créteil Cedex France; G. C. MEL-LAU, Justus-Liebig - Universität, Physikalisch-Chemisches Institut, Heinrich-Buff-Ring 58, D-35392 Gießen, Germany.

Using high-resolution Fourier transform spectra of mono deuterated formaldehyde (HDCO) recorded in the 5.8 μm spectral range at Giessen (Germany), we carried out an extensive analysis of the strong ν_{3} fundamental band (carbonyl stretching mode) at 1724.2676 cm^{−1}, starting from results of a previous analysis ^{a}. For this hybrid band (with both A- and B-type transitions) the analysis was pursued up to high rotational quantum numbers. In this way, it was possible to evidence resonances which perturb the ν_{3} lines which are due to the existence of the 2ν_{5} (at 2059 cm^{−1}) and ν_{5}+ν_{6} (at 2087 cm^{−1}) dark bands ^{b}. In addition a local resonance is perturbing the 3^{1} levels which is due to a crossing with the 4^{1} energy levels. However the 4^{1} state is also involved in strong vibration-rotation interactions coupling the {5^{1},6^{1},4^{1}} system of resonating states of HDCO^{c}. Therefore the final energy levels calculation which was performed for the {5^{1},6^{1},4^{1},3^{1},5^{2},5^{1}6^{1}} resonating states accounts for the observed A- type, B- type C-type Coriolis (and/or) Fermi resonances. In this way it was possible to reproduce the observed line positions, within their experimental uncertainties. Finally using a ν_{3} band intensity available in the literature ^{d} we generated, for the first time, a list of line parameters (positions and intensities) for the 5.8 μm band of HDCO.

aJohns JWC, McKellar ARW., J Mol Spectrosc 1977; 64: 327-339

bν_{4}, ν_{5} and ν_{6} correspond to the CHD bending (at 1396 cm^{−1}), the CHD rocking (at 1028 cm^{−1}) and the CHD out of plane (at 1059 cm^{−1}) modes, respectively

cA. Perrin, J. M. Flaud, L. Margulès, J. Demaison, H. Mäder and S. Wörmke, J. Mol. Spectrosc. 216, 214 (2002)

dGratien, Nilsson, Doussin, Johnson, Nielsen, Stenstrom and Picquet-Varrault, J. Phys. Chem. 111, 11506 (2007)